OpenStudy (anonymous):

Please help!! What are the restrictions for x(x-1)(x-2)? I already solved the problem and this is the denominator...I'm not sure how to find the restrictions.

OpenStudy (anonymous):

if this is denominator, it can't be 0. So find it's roots : x(x-1)(x-2)=0 It will be at x=0, x=1, and x=2

OpenStudy (anonymous):

Oh ok. So do I set the whole thing equal to 0 or x(x-1)=0 or x(x-2)=0? The restircitons are 0, 1, and 2?

OpenStudy (anonymous):

@myko ??? Wait so what are the restrictions??

OpenStudy (anonymous):

The hall thing. The restrictions are the roots.

OpenStudy (anonymous):

Oh ok! So it's 0, 1, and 2?

OpenStudy (anonymous):

as you know, the fraction can't have 0 denominator, because then it's not defined. Sincé x(x-1)(x-2) is the denominator, any value of x, that makes it 0 is not alowed. For example, when x=2, we eill have 2(2-1)(2-2)=2(1)(0)=0. So x=2 is not alowed. It is a restriction. Same for x=0, and x=1

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

THANK YOU!!

OpenStudy (anonymous):

yw